OpenStudy (australopithecus):

How can I use Demorgins law to express: AB'C'D' + ACB' + CD' = Y In terms of NAND Operators

ganeshie8 (ganeshie8):

you can represent it using 4 NAND gates after using demorgan laws

OpenStudy (australopithecus):

So Demorgins law is X + Y = XY and (XY)' = x' + y' I Expand it to A + B' + C' + D' + A + C + B' + C + D' = (A + A) + (B' + B') + (C + C') + (D' + D') = A + B' + C' + D' What am I doing wrong? I think I might just be confused about how to properly utilize the law I'm aware a NAND Gate is just a Not And Gate

ganeshie8 (ganeshie8):

Demorgan laws :- (X + Y)' = X'Y' and (XY)' = x' + y'

OpenStudy (australopithecus):

So, X + Y =/= XY

ganeshie8 (ganeshie8):

yes. they are not same.

ganeshie8 (ganeshie8):

X+Y is OR gate XY is AND gate

OpenStudy (australopithecus):

My Profs notes are kind of messed up, so my list of rules are pretty sketchy

OpenStudy (australopithecus):

yeah that makes sense blah must not be thinking

ganeshie8 (ganeshie8):

ok look at the two demorgan laws again, before we begin

ganeshie8 (ganeshie8):

you cannot expand or do anything with the given expression as it is.

ganeshie8 (ganeshie8):

Given function : Y = AB'C'D' + ACB' + CD'

OpenStudy (australopithecus):

So I have to use boolean algebra to manipulate it right?

ganeshie8 (ganeshie8):

Ima show you how to sovle..

OpenStudy (australopithecus):

Ok thank you, just learned this stuff an hour ago and any help is appreciated greatly

ganeshie8 (ganeshie8):

\(Y = AB'C'D' + ACB' + CD' \) \(Y = \overline{\overline{AB'C'D' + ACB' + CD'}} \)

ganeshie8 (ganeshie8):

you understand the bars right ?

OpenStudy (australopithecus):

Yes bars just mean the inverse, my latex isnt working give me a second going to restart browser

ganeshie8 (ganeshie8):

bar and a tick mean the same thing : NOT

ganeshie8 (ganeshie8):

oh ok.. take ur tiem :)

OpenStudy (australopithecus):

ha that statement is confusing, well not really but you know

OpenStudy (australopithecus):

ok back latex is working for me

ganeshie8 (ganeshie8):

yea.. NOT is a NOT gate

ganeshie8 (ganeshie8):

ok lets continue

OpenStudy (australopithecus):

Alright

OpenStudy (australopithecus):

thanks again for the help

ganeshie8 (ganeshie8):

\(Y = AB'C'D' + ACB' + CD' \) \(Y = \overline{\overline{AB'C'D' + ACB' + CD'}}\) \(Y = \overline{\overline{\color{green}{AB'C'D'} + \color{purple}{ACB' + CD'}}}\)

ganeshie8 (ganeshie8):

apply demorgan for green and purple terms

OpenStudy (australopithecus):

How, I'm sorry I'm still confused on how to manipulate these can you give me an example?

ganeshie8 (ganeshie8):

\(Y = AB'C'D' + ACB' + CD' \) \(Y = \overline{\overline{AB'C'D' + ACB' + CD'}}\) \(Y = \overline{\overline{\color{green}{AB'C'D'} + \color{purple}{ACB' + CD'}}}\) \(Y = \overline{ \overline{\color{green}{(AB'C'D')}} . \overline{ \color{purple}{(ACB' + CD')}} } \)

ganeshie8 (ganeshie8):

have a look i used this :- \(\overline{\color{green}{X}+\color{purple}{Y}} = \overline{\color{green}{X}} . \overline{\color{purple}{Y}\)

OpenStudy (australopithecus):

Oh Oh I see, so ( (AB'C'D')(ACB')(CD') )')' how does this bring us any close to the solution

ganeshie8 (ganeshie8):

contune and see hw we get all NAND gates

ganeshie8 (ganeshie8):

\(Y = AB'C'D' + ACB' + CD' \) \(Y = \overline{\overline{AB'C'D' + ACB' + CD'}}\) \(Y = \overline{\overline{\color{green}{AB'C'D'} + \color{purple}{ACB' + CD'}}}\) \(Y = \overline{ \overline{\color{green}{(AB'C'D')}} . \overline{ \color{purple}{(ACB' + CD')}} } \) \(Y = \overline{ \overline{\color{green}{AB'C'D'}} . \color{purple}{\overline{ACB'} . \overline{CD'}} } \)

ganeshie8 (ganeshie8):

we're done

OpenStudy (australopithecus):

oh I see you just rearrange the variables

OpenStudy (australopithecus):

and you use double inverse because ((x)')' Thank you so much

ganeshie8 (ganeshie8):

yes i think u see now, those are 3 NAND gates for minterms and one more outer NAND gate combining those

OpenStudy (australopithecus):

:) you have been so much help this problem has been driving me crazy :)

ganeshie8 (ganeshie8):

np :) digital electronics is easy to learn... 2-3 days more im sure u wil feel confident :)